# From Pixels to Insights: A Survey on Automatic Chart Understanding in the Era of Large Foundation Models

### v2

@@ Abstract
Data visualization in the form of charts plays a pivotal role in data analysis, offering critical insights and aiding in informed decision-making. Automatic chart understanding has witnessed significant advancements with the rise of large foundation models in recent years. Foundation models, such as large language models, have revolutionized various natural language processing tasks and are increasingly being applied to chart understanding tasks. This survey paper provides a comprehensive overview of the recent developments, challenges, and future directions in chart understanding within the context of these foundation models. We review fundamental building blocks crucial for studying chart understanding tasks. Additionally, we explore various tasks and their evaluation metrics and sources of both charts and textual inputs. Various modeling strategies are then examined, encompassing both classification-based and generation-based approaches, along with tool augmentation techniques that enhance chart understanding performance. Furthermore, we discuss the state-of-the-art performance of each task and discuss how we can improve the performance. Challenges and future directions are addressed, highlighting the importance of several topics, such as domain-specific charts, lack of efforts in developing evaluation metrics, and agent-oriented settings. This survey paper serves as a comprehensive resource for researchers and practitioners in the fields of natural language processing, computer vision, and data analysis, providing valuable insights and directions for future research in chart understanding leveraging large foundation models. The studies mentioned in this paper, along with emerging new research, will be continually updated at: https://github.com/khuangaf/Awesome-Chart-Understanding.
@@ 1 Introduction
Challenges and Opportunities in Chart Understanding Amidst the Era of Large Foundation Models  Traditional chart understanding work [15, 7] focuses on finetuning methods that generally experience limitations with respect to domain portability and reasoning robustness. Excitingly, the advent of large vision-language foundation models (e.g., GPT-4V [16], LLaVA [17]) has spawned a paradigm shift in automated reasoning capabilities, catalyzing unprecedented advances across various multimedia cognitive tasks – including progress in strong zero/few-shot reasoning capabilities through text-based prompting. Despite the impressive improvements, the domain of chart understanding still presents major challenges. Charts present a unique set of obstacles owing to their rich visual representations and nuanced semantics, delivered through complex and creative organization of information manifested in text, numerical data, plots, and graphics objects. From bar charts and line graphs to pie charts and scatter plots, each chart type employs a distinct visual syntax to convey data relationships, requiring sophisticated interpretative mechanisms beyond mere pixel-level pattern recognition. Charts serve as a conduit for unearthing profound insights such as emerging trends, outliers that challenge assumptions, and relationships among variables that may not be immediately apparent from raw data in tabular form alone. They enable comparative analyses across data points, offering a visual platform for juxtaposing varied entities or time periods succinctly. Furthermore, the intrinsic diversity of underlying datasets, ranging from simple numerical relations to intricate multidimensional entities, adds another layer of complexity to the chart understanding task. Despite these challenges, automated chart understanding offers a gateway to unlock actionable insights buried within the pixels of visual narratives. By harnessing the capabilities of large foundation models, chart understanding demonstrates improved potential in bridging the gap between raw visual data and meaningful insights, thereby enabling high-level information extraction, logical reasoning, and decision-making.
@@
We first define automatic chart understanding and the fundamental building blocks to the problem formulation in Section 2. We discuss the multifaceted nature of chart understanding, encompassing tasks from interpreting chart visuals to analyzing underlying data. We also illustrate how chart understanding stands amidst the intersection of related works in natural image understanding [24], table understanding [25], and document understanding [26], as well as outling the structural modeling components integral to chart understanding, such as vision encoders, OCR modules, and text decoders, and their role in transforming raw chart images and text queries into meaningful insights. Then, in Section 3, we examine the datasets driving chart understanding research and the metrics for model evaluation. This section analyzes the sources, diversity, and limitations of these datasets, providing insights into the current chart understanding data landscape. It also reviews various evaluation metrics, underlining the necessity for robust and nuanced assessment methods. With insights from these characteristics, we further provide trending popular modeling strategies for automatic chart understanding. Section 4 delves into the diverse modeling strategies in chart understanding, including adaptations from natural image understanding, vision-language pre-training, and foundation models like large language models (LLMs) and large vision-language models (LVLMs). In particular, we emphasize the impact of choices in vision encoders and text decoders on model effectiveness and discuss tool augmentation’s role in chart understanding. We conclude this section by showcasing the state-of-the-art performance on different chart understanding tasks and how we can improve upon them.
@@ 2.2.1 Natural Image Understanding
Natural image understanding focuses on the interpretation and analysis of photographic images, which capture scenes from the real world. These images can include landscapes, people, objects, and everyday situations. The objective of this field is to enable machines to recognize and make sense of visual information in a manner that replicates human vision. Natural image understanding encompasses tasks such as image captioning [27, 28] and visual question answering [29, 30]. The core difference between natural image understanding and chart understanding lies in the type of visual content and the nature of the interpretation tasks involved. While natural image understanding deals with unstructured visual data that represents natural scenes, chart understanding focuses on structured visual representations of data, such as bar graphs, line charts, and pie charts. Charts are designed to communicate specific quantitative information and trends, making the interpretation task centered around data extraction, recognition of graphical elements, and comprehension of the data narrative conveyed by the chart.
@@ 2.2.2 Table Understanding
Furthermore, while it might be arguable that chart understanding tasks can be simplified into table understanding tasks by converting charts into underlying data tables. data tables are infrequently published alongside charts, and the process of automatically extracting data tables is fraught with challenges. The diversity of chart types and visual styles, coupled with the intricacies of reasoning about visual occlusion, makes automatic data table extraction prone to errors. This complexity underscores the distinct and intricate nature of chart understanding, highlighting its unique challenges compared to table understanding.
@@ 2.3 Problem Formulation and Fundamental Building Blocks
Vision Encoder: It is often essential to extract visual features from the input chart image in order to understand the relations and spatial arrangements between the graphical elements and scene text in the chart. Therefore, conventional chart understanding models usually employ a vision encoder $\textsc{Encoder}_{\text{vis}}$ to map the input chart image to a visual feature matrix: $\mathbf{H}_{c}=\textsc{Encoder}_{\text{vis}}(\mathbf{c})$. Recently developed models based on Large Vision-Language Models (LVLMs), such as ChartLlama [38] and ChartAssistant [39], employ an additional projection layer $\textsc{Projector}_{\text{vt}}$ to better align text and visual representations: $\mathbf{H}_{c}=\textsc{Projector}_{\text{vt}}(\textsc{Encoder}_{\text{vis}}(\mathbf{c}))$.
Chart-to-Table Conversion Module: The underlying data table $\mathbf{T}$ of an input chart provides a fully structured textual representation of its raw data, which can help language models better understand the information presented in the input chart. In real-world applications, the underlying data tables of charts may not readily accessible. Therefore, various chart understanding approaches employ a chart-to-table translation module to extract a data table $\hat{\mathbf{T}}$ from the input chart. An extracted table can also be linearized into a sequence of table tokens $[\hat{\mathbf{T}}_{1},\ldots,\hat{\mathbf{T}}_{l^{\hat{\mathbf{T}}}}]$, where $l^{\hat{\mathbf{T}}}$ denotes the number of elements in $\hat{\mathbf{T}}$.
@@ 3.2.1 Source of Charts
The debate between the use of synthetic versus real-world charts is pivotal in the chart understanding field. Synthetic charts, as generated by visualization tools according to predefined rules, offer the advantage of full control over the dataset, including the variety of chart types, the range and distribution of values, and the incorporation of specific challenges to test model robustness. This controlled setting allows for the systematic study of model behaviors and the identification of specific weaknesses. Furthermore, the process of generating synthetic charts is scalable and can easily produce large volumes of data necessary for training sophisticated machine learning models. However, the primary drawback of synthetic charts is their potential lack of realism. The generated charts may not fully capture the complexity and noise found in real-world charts, which can lead to poor performance in real-life scenarios for models trained on synthetic charts. Moreover, another challenge for automatically creating synthetic is balancing the desire for visual diversity and the need for the charts to remain comprehensible. We often aim to introduce variation in the styles of the charts, such as through randomized colors or plot types, to create a dataset that encompasses a wide range of visual appearances. However, ensuring that the resulting charts make visual sense poses a considerable difficulty. For instance, in a bar chart designed to compare different entities, if the bars corresponding to these entities are colored very similarly due to randomization, it may lead to confusion and impair the chart’s readability. Such issues underscore the complexity of automatically generating synthetic charts that are not only diverse but also logically coherent and visually clear.
@@
Conversely, training on real-world charts could potentially lead to models better attuned to the nuances of real-world data visualizations, improving their performance on tasks outside controlled experimental settings. However, the use of real-world charts comes with its own set of challenges. First, collecting a large and diverse enough dataset to train robust models can be difficult and time-consuming. Additionally, the available charts may be biased towards certain domains, styles, or visual layouts, depending on the sources. Furthermore, labeling real-world charts for tasks such as question answering requires significant expert effort, making the dataset creation process slower and more expensive.
@@ 3.2.3 Plots
While bar, line, and pie charts form the core of most datasets, some recent additions have begun to incorporate area plots, which are valuable for showing cumulative quantities and changes over time. Notably, OpenCQA [5] and Chart-to-Text [8] include area plots, thus expanding the scope of analysis possible with these datasets. Additionally, they also include scatter plots, which are effective for identifying correlations and distributions of data points.
@@
Plot Types Dataset Line Bar Area Pie Scatter Others FigureQA [2] DVQA [1] LEAF-QA [43] LEAF-QA++ [44] PlotQA [4] ChartQA [3] MapQA [45] PaperQA [46] SciGraphQA [47] MMC-Bench [48] ChartBench [49] ArXivQA [50] OpenCQA [5] FigCAP [51] Chart2text [52] SciCap [6] LineCap [53] ChaTa+ [54] Chart-to-Text [8] ChartSumm [55] VisText [9] FigCaps-HF [56] ArXivCap [50] ChartFC [13] ChartCheck [12] Chocolate [14]
@@ 3.3 Evaluation Metric
Note that this setting resembles factual inconsistency detection [14] and fact-checking [12] but is distinct in its use of contrasting prompts and the framing of textual queries as question-answer pairs instead of declarative statements.
Chart-to-Table Conversion. Two metrics have been developed for the chart-to-table conversion task, Relative Number Set Similarity (RNSS) [3] and Relative Mapping Similarity (RMS) [10]. For this task, we can view tables as unordered collections of mappings from a row-and-column header $(r,c)$ to a single value $v$. RNSS represents each entry of the predicted table using the values only. Specifically, $\{\hat{\mathbf{T}}^{v}_{i}\}_{1\leq i\leq N}$ and $\{\mathbf{T}^{v}_{j}\}_{1\leq j\leq M}$ represent the set of values in the ground truth and generated tables, respectively. RNSS defines the similarity $S$ between each pair of values:
@@
Long-form Chart Understanding Tasks. The tasks discussed above are relatively easy to evaluate since the ground truths are either short answers or long answers that can be easily decomposed into short answers due to the structural nature of the outputs. Evaluating tasks that involve long-form texts, including Open-ended Chart Question Answering, Chart Captioning, and Chart Caption Factual Error Correction is more challenging since the outputs are much longer and cannot be easily broken down into short sub-outputs. Most of the prior work use reference-based evaluation metrics, which score the outputs against reference captions or answers. Some adopt lexical-based metrics, such as ROUGE [60] and BLEU [61], while others use Large Language Models (LLMs) as evaluation metrics [38], motivated by the text generation fields [62, 63]. However, these metrics have several drawbacks. First, these metrics do not correlate well with faithfulness, or factual consistency, between the chart and the outputs, as demonstrated by prior work in the field of summarization [64, 65, 66]. Second, they depend on the quality (e.g. coverage) of the reference outputs. Taking chart captioning as an example, the majority of instances in existing datasets do not cover all key insights from the chart. Therefore, it is possible that the generated caption is faithful to the given charts but has a low lexical overlap with the reference caption. To address these issues, ChartVE [14] formulates the factual inconsistency detection problem as a visual entailment task. ChartVE learns to predict the entailment probability $E(\mathbf{c},\mathbf{q})$ from the input chart $\mathbf{c}$ to a caption sentence $\mathbf{q}$. The ChartVE score is computed by taking the minimum of all caption sentences $\mathbf{q}\in\mathbf{Q}$:

### v4

@@ Abstract
Data visualization in the form of charts plays a pivotal role in data analysis, offering critical insights and aiding in informed decision-making. Automatic chart understanding has witnessed significant advancements with the rise of large foundation models in recent years. Foundation models, such as large language models, have revolutionized various natural language processing tasks and are increasingly being applied to chart understanding tasks. This survey paper provides a comprehensive overview of the recent developments, challenges, and future directions in chart understanding within the context of these foundation models. We review fundamental building blocks crucial for studying chart understanding tasks. Additionally, we explore various tasks and their evaluation metrics and sources of both charts and textual inputs. Various modeling strategies are then examined, encompassing both classification-based and generation-based approaches, along with tool augmentation techniques that enhance chart understanding performance. Furthermore, we discuss the state-of-the-art performance of each task and discuss how we can improve the performance. Challenges and future directions are addressed, highlighting the importance of several topics, such as domain-specific charts, lack of efforts in developing evaluation metrics, and agent-oriented settings. This survey paper aims to provide valuable insights and directions for future research in chart understanding leveraging large foundation models.
@@ 1 Introduction
Challenges and Opportunities in Chart Understanding Amidst the Era of Large Foundation Models  Traditional chart understanding work [15, 7] focuses on finetuning methods that generally experience limitations with respect to domain portability and reasoning robustness. Excitingly, the advent of large vision-language foundation models (e.g., GPT-4V [16], LLaVA [17]) has spawned a paradigm shift in automated reasoning capabilities, catalyzing unprecedented advances across various multimedia cognitive tasks – including progress in strong zero/few-shot reasoning capabilities through text-based prompting. Despite the impressive improvements, the domain of chart understanding still presents major challenges. Charts present a unique set of obstacles owing to their rich visual representations and nuanced semantics, delivered through complex and creative organization of information manifested in text, numerical data, plots, and graphics objects. From bar charts and line graphs to pie charts and scatter plots, each chart type employs a distinct visual syntax to convey data relationships, requiring sophisticated interpretative mechanisms beyond mere pixel-level pattern recognition. Charts serve as a means to uncover important insights, such as emerging trends, outliers that challenge assumptions, and relationships among variables that are not immediately apparent from raw data in tabular form. They enable comparative analyses across data points, providing a visual platform for summarizing varied entities or time periods clearly. Furthermore, the intrinsic diversity of underlying datasets, ranging from simple numerical relations to intricate multidimensional entities, adds another layer of complexity to the chart understanding task. Despite these challenges, automated chart understanding offers a gateway to unlock actionable insights buried within the pixels of visual narratives. By harnessing the capabilities of large foundation models, chart understanding demonstrates improved potential in bridging the gap between raw visual data and meaningful insights, thereby enabling high-level information extraction, logical reasoning, and decision-making.
@@
We first define automatic chart understanding and the fundamental building blocks to the problem formulation in Section 2. We discuss the multifaceted nature of chart understanding, encompassing tasks from interpreting chart visuals to analyzing underlying data. We also illustrate how chart understanding overlaps with related works in natural image understanding [24], table understanding [25], and document understanding [26]. We explain the essential modeling components for chart understanding, such as vision encoders, OCR modules, and text decoders, and their role in converting raw chart images and text queries into useful insights.Then, in Section 3, we examine the datasets driving chart understanding research and the metrics for model evaluation. This section analyzes the sources, diversity, and limitations of these datasets, providing insights into the current chart understanding data landscape. It also reviews various evaluation metrics, underlining the necessity for robust and nuanced assessment methods. With insights from these characteristics, we further provide trending popular modeling strategies for automatic chart understanding. Section 4 delves into the diverse modeling strategies in chart understanding, including adaptations from natural image understanding, vision-language pre-training, and foundation models like large language models (LLMs) and large vision-language models (LVLMs). In particular, we emphasize the impact of choices in vision encoders and text decoders on model effectiveness and discuss tool augmentation’s role in chart understanding. We conclude this section by showcasing the state-of-the-art performance on different chart understanding tasks and how we can improve upon them.
@@ 2.2.1 Natural Image Understanding
Natural image understanding focuses on the interpretation and analysis of photographic images, which capture scenes from the real world. These images can include landscapes, people, objects, and everyday situations. The objective of this field is to enable machines to recognize and make sense of visual information in a manner that replicates human vision. Natural image understanding encompasses tasks such as image captioning [27, 28] and visual question answering [29, 30]. The core difference between natural image understanding and chart understanding lies in the type of visual content and the nature of the interpretation tasks involved. While natural image understanding deals with unstructured visual data that represents natural scenes, chart understanding focuses on structured visual representations of data, such as bar graphs, line charts, and pie charts. Charts are designed to communicate specific quantitative information and trends. This makes the interpretation task centered around data extraction, recognition of graphical elements, and comprehension of the data narrative conveyed by the chart.
@@ 2.2.2 Table Understanding
Furthermore, while it might be arguable that chart understanding tasks can be simplified into table understanding tasks by converting charts into underlying data tables. However, data tables are infrequently published alongside charts, and the process of automatically extracting data tables is fraught with challenges. The diversity of chart types and visual styles, coupled with the intricacies of reasoning about visual occlusion, makes automatic data table extraction prone to errors. This complexity underscores the distinct and intricate nature of chart understanding, highlighting its unique challenges compared to table understanding.
@@ 2.3 Problem Formulation and Fundamental Building Blocks
Vision Encoder: It is often essential to extract visual features from the input chart image in order to understand the relations and spatial arrangements between the graphical elements and scene text in the chart. Therefore, conventional chart understanding models usually employ a vision encoder $\textsc{Encoder}_{\text{vis}}$ to map the input chart image to a visual feature matrix: $\mathbf{H}_{c}=\textsc{Encoder}_{\text{vis}}(\mathbf{c})$.
Chart-to-Table Conversion Module: The underlying data table $\mathbf{T}$ of an input chart provides a fully structured textual representation of its raw data. This can help language models better understand the information presented in the input chart. In real-world applications, the underlying data tables of charts may not readily accessible. Therefore, various chart understanding approaches employ a chart-to-table translation module to extract a data table $\hat{\mathbf{T}}$ from the input chart. An extracted table can also be linearized into a sequence of table tokens $[\hat{\mathbf{T}}_{1},\ldots,\hat{\mathbf{T}}_{l^{\hat{\mathbf{T}}}}]$, where $l^{\hat{\mathbf{T}}}$ denotes the number of elements in $\hat{\mathbf{T}}$.
@@
Additionally, there are some variations in the way that input charts are encoded. For example, recently developed Large Vision-Language Models (LVLMs) like ChartLlama [41] and ChartAssistant [42] employ an additional projection layer $\textsc{Projector}_{\text{vt}}$ to better align text and visual representations: $\mathbf{H}_{c}=\textsc{Projector}_{\text{vt}}(\textsc{Encoder}_{\text{vis}}(\mathbf{c}))$.
@@ 3.1 Tasks
Due to the development of LVLMs, recent studies have introduced more challenging tasks. For example, tasks such as Chart Redrawing and Chart Referring QA have been proposed to push the boundaries of what these models can achieve [58, 42]. While these emerging tasks present exciting avenues for future research, they are still in their nascent stages and thus not comprehensively covered in this survey.
@@ 3.2.1 Source of Charts
Additionally, with the advancement of LLMs, there is a growing trend of utilizing LLMs to produce synthetic charts. Examples include ChartLlama [41], ChartX [58], and SimChart9K [59], which simulate data across various topics and generate code to plot charts using LLMs.
The debate between the use of synthetic versus real-world charts is crucial in the chart understanding field. Synthetic charts, as generated by visualization tools according to predefined rules, offer the advantage of full control over the dataset, including the variety of chart types, the range and distribution of values, and the incorporation of specific challenges to test model robustness. This controlled setting allows for the systematic study of model behaviors and the identification of specific weaknesses. Furthermore, the process of generating synthetic charts is scalable and can easily produce large volumes of data necessary for training sophisticated machine learning models. However, the primary drawback of synthetic charts is their potential lack of realism. The generated charts may not fully capture the complexity and noise found in real-world charts, which can lead to poor performance in real-life scenarios for models trained on synthetic charts. Moreover, another challenge for automatically creating synthetic is balancing the desire for visual diversity and the need for the charts to remain comprehensible. We often aim to introduce variation in the styles of the charts, such as through randomized colors or plot types, to create a dataset that encompasses a wide range of visual appearances. However, ensuring that the resulting charts make visual sense poses a considerable difficulty. For instance, in a bar chart designed to compare different entities, if the bars corresponding to these entities are colored very similarly due to randomization, it may lead to confusion and impair the chart’s readability. Such issues underscore the complexity of automatically generating synthetic charts that are not only diverse but also logically coherent and visually clear.
@@
Conversely, training on real-world charts could potentially lead to models better attuned to the nuances of real-world data visualizations. This could improve their performance on tasks outside controlled experimental settings. However, the use of real-world charts comes with its own set of challenges. First, collecting a large and diverse enough dataset to train robust models can be difficult and time-consuming. Additionally, the available charts may be biased towards certain domains, styles, or visual layouts, depending on the sources. Furthermore, labeling real-world charts for tasks such as question answering requires significant expert effort, making the dataset creation process slower and more expensive.
@@ 3.2.2 Source of Textual Inputs
More recently, due to the advancement of text generation, work like ChartX [58] and SimChart9K [60] use LLMs to generate diverse and high-quality textual inputs.
@@ 3.2.3 Plots
While bar, line, and pie charts form the core of most datasets, some recent additions have begun to incorporate area plots, which are valuable for showing cumulative quantities and changes over time. Notably, OpenCQA [5] and Chart-to-Text [8] include area plots, thus expanding the scope of analysis possible with these datasets. Additionally, ChartX [58] is currently the most comprehensive chart understanding dataset, encompassing a total of 18 chart types ranging from bar charts to candlestick plots.
@@
Plot Types Dataset Line Bar Area Pie Scatter Others FigureQA [2] DVQA [1] LEAF-QA [43] LEAF-QA++ [44] PlotQA [4] ChartQA [3] MapQA [45] PaperQA [46] SciGraphQA [47] MMC-Bench [48] ChartBench [49] ArXivQA [50] OpenCQA [5] FigCAP [51] Chart2text [52] SciCap [6] LineCap [53] ChaTa+ [54] Chart-to-Text [8] ChartSumm [55] VisText [9] FigCaps-HF [56] ArXivCap [50] ChartFC [13] ChartCheck [12] Chocolate [14] ChartLlama [41] ChartX [58]
@@ 3.3 Evaluation Metric
Note that this setting resembles factual inconsistency detection [14] and fact-checking [12]. However, it is distinct in its use of contrasting prompts and the framing of textual queries as question-answer pairs instead of declarative statements.
Chart-to-Table Conversion. Three metrics have been developed for the chart-to-table conversion task, Relative Number Set Similarity (RNSS) [3], Relative Mapping Similarity (RMS) [10], and Structuring Chart-oriented Representation Metric (SCRM) [60]. For this task, we can view tables as unordered collections of mappings from a row-and-column header $(r,c)$ to a single value $v$. RNSS represents each entry of the predicted table using the values only. Specifically, $\{\hat{\mathbf{T}}^{v}_{i}\}_{1\leq i\leq N}$ and $\{\mathbf{T}^{v}_{j}\}_{1\leq j\leq M}$ represent the set of values in the ground truth and generated tables, respectively. RNSS defines the similarity $S$ between each pair of values:
@@
Unlike RNSS and RMS, SCRM represents each table entry as a triplet through the Structured Triplet Representations (STR) framework. STR is designed to efficiently and robustly represent the complex data relations within chart information by capturing the relationships between row and column headers and their corresponding values. SCRM is computed by first calculating entity and value matching metrics. Thus, the SCRM metric provides a comprehensive evaluation metric to assess the performance of chart perception from structured information extraction, addressing the inherent complexities associated with data relations in visual charts.
@@
Long-form Chart Understanding Tasks. The tasks discussed above are relatively easy to evaluate since the ground truths are either short answers or long answers that can be easily decomposed into short answers due to the structural nature of the outputs. Evaluating tasks that involve long-form texts, including Open-ended Chart Question Answering, Chart Captioning, and Chart Caption Factual Error Correction is more challenging since the outputs are much longer and cannot be easily broken down into short sub-outputs. Most of the prior work use reference-based evaluation metrics, which score the outputs against reference captions or answers. Some adopt lexical-based metrics, such as ROUGE [63], BLEU [64], and CIDEr [65], another line of studies use semantic-focused metrics like perplexity and BERTScore [66], and BLEURT [67], others use Large Language Models (LLMs) as evaluation metrics [41], motivated by the text generation fields [68, 69]. However, these metrics have several drawbacks. First, these metrics do not correlate well with faithfulness, or factual consistency, between the chart and the outputs, as demonstrated by prior work in the field of summarization [70, 71, 72]. Second, they depend on the quality (e.g. coverage) of the reference outputs. Taking chart captioning as an example, the majority of instances in existing datasets do not cover all key insights from the chart. Therefore, it is possible that the generated caption is faithful to the given charts but has a low lexical overlap with the reference caption. To address these issues, ChartVE [14] formulates the factual inconsistency detection problem as a visual entailment task. ChartVE learns to predict the entailment probability $E(\mathbf{c},\mathbf{q})$ from the input chart $\mathbf{c}$ to a caption sentence $\mathbf{q}$. The ChartVE score is computed by taking the minimum of all caption sentences $\mathbf{q}\in\mathbf{Q}$: